OpenStudy (anonymous):

What is 3/2 times -5

OpenStudy (anonymous):

Actually i'm not sure, you could divide the denominator by the numerator and multiply by -5

OpenStudy (anonymous):

-5 is equal to \[\frac{ -5 }{ 1 }\] So, do \[\frac{ 3 }{ 2 } * \frac{-5} {1}\] To multiply fractions, you just multiply the numerators and denominators straight across, so you would get \[\frac{ 3 * -5 }{ 2 * 1 }\] Simplify: \[\frac{ -15 }{ 2 }\] Hope I helped! If you have any more questions, tag me or message me, and I'll do my best to help out :D